# Multidimensional empirical wavelet transform

### v1

@@ Abstract
The empirical wavelet transform is a data-driven time-scale representation consisting of adaptive filters. Its robustness to data has made it the subject of intense developments and an increasing number of applications in the last decade. However, it has been mostly studied theoretically for signals so far and its extension to images is limited to a particular mother wavelet. This work presents a general framework for multidimensional empirical wavelet transform from any mother wavelet. In addition, it provides conditions to build wavelet frames for both continuous and discrete transforms.
@@ 1 Introduction
The wavelet transform is a reference tool for time-scale representation used in many signal and image processing techniques, such as denoising, deconvolution and texture segmentation. Originally, it consists of projecting data onto wavelet filters that are built from a mother wavelet which is scaled and modulated independently of the data. In practice, this leads to the construction of wavelet filters based on a prescribed scheme, such as a dyadic decomposition. Although this approach is widely used in contemporary research, it is not guaranteed to be optimal for the data at hand, since a prescribed scheme does not take into account the specificity of the underlying Fourier spectrum. Therefore, data-driven filtering approaches have received much attention to provide an accurate time-scale representation that is robust to the data. Among them, inspired by the empirical mode decomposition [13], the empirical wavelet transform, introduced in [6], has been the subject of an increasing interest in the last decade through a continuous development and numerous applications in various fields, as reviewed in [14, 21]. To name a few applications, we can mention seismic time-frequency analysis [20], electrocardiogram (ECG) signal compression [17], epileptic seizure detection [1, 25], speech recognition [19], time series forecasting [23], glaucoma detection [22], hyperspectral image classification [24], texture segmentation [15], multimodal medical image fusion [26] and cancer histopathological image classification [5].
@@
The construction of empirical wavelet systems consists of two steps: $(i)$ extracting supports of the harmonic modes of the function under study, and $(ii)$ constructing empirical wavelet filters that are mostly (compactly or very rapidly decaying) supported in the Fourier domain by the extracted supports. This construction is the core of an ongoing and active research effort for one-dimensional (1D) and two-dimensional (2D) functions. For 1D functions, the detection of the segments supporting each harmonic modes in the Fourier domain is usually performed by extracting the lowest minima between them using a scale-space representation [10]. For 2D functions, several different techniques have been proposed to delimit the supports of the harmonic modes, such as the Curvelet [11], Watershed [16] and Voronoi [8] partitioning methods. The construction of wavelet filters based on the detected supports is usually done in the Fourier domain. In the 1D domain, [7] proposed a general framework to build continuous empirical wavelet filters from a mother wavelet, such as the Meyer, Shannon or Gabor wavelets. These 1D empirical wavelet systems are written as modulations and dilations of a mother wavelet based on the segments that divide the Fourier line. Such systems have been shown to induce both continuous and discrete frames in [9]. In the 2D domain, empirical wavelet filters have been designed following the Littlewood-Palley wavelet formulation for various Fourier partitioning methods [8, 11, 16]. However, the proposed construction is only valid for discrete functions and is entirely based on the distance to the support boundaries, which limits its extension to classic mother wavelets. So far, no construction of empirical wavelets in higher dimensions has been proposed. In addition, empirical wavelet filters for real-valued functions are built from supports that take into account the symmetry of the corresponding Fourier transform, but these have been little studied theoretically.
@@
This work aims to provide a general framework for empirical wavelet transforms of multidimensional functions, thus extending the 1D framework in [7, 9]. We show that we can build empirical wavelets on Fourier supports, symmetric or not, from any mother wavelet defined in the Fourier domain, using diffeomorphisms that map these supports to the mother wavelet’s Fourier support. Both continuous and discrete transforms are considered. In addition, conditions for the construction of wavelet frames are examined.
@@
The paper is organized as follows. The construction of the multidimensional empirical wavelet systems and the resulting transforms is described in Section 3. Theoretical results on these systems as frames are given in Section 4. The special case of Fourier supports resulting from affine deformations of the mother wavelet’s Fourier support is studied in Section 5. Finally, in Section 6, specific 2D wavelet systems are tailored from classic mother wavelets and studied numerically on images. A Matlab toolbox will be made publicly available at the time of publication.
@@ 2 Notations and reminders
We respectively denote $\partial\Omega$ and $\overline{\Omega}$ the boundary and closure of a set $\Omega\in{\mathbb{R}}^{N}$. We denote $\Upsilon^{+}=\{n\in\Upsilon\mid n\geq 0\}$ the subset of positive elements of a set $\Upsilon\in{\mathbb{Z}}$. The Jacobian of a differentiable function $\varphi$ is denoted $J_{\varphi}$. We recall that a function $\varphi$ is called a diffeomorphism if it is infinitely differentiable and invertible of inverse infinitely differentiable.
@@
with $({\mathcal{V}},\mu)$ a measure set. In particular, $\{g_{p}\}_{p\in{\mathcal{V}}}$ is called a tight frame if $A_{1}=A_{2}$ and a Parseval frame if $A_{1}=A_{2}=1$. Morover, it is well known that $\{g_{p}\}_{p\in{\mathcal{V}}}$ is a tight frame if and only if there exist $A>0$ such that, for every $f\in{\mathrm{L}^{2}}({\mathbb{R}}^{N})$,
$f(x)=\frac{1}{A}\int_{{\mathcal{V}}}\langle f,g_{p}\rangle\,g_{p}(x)\,{\mathrm{d}}\mu(p).$
A frame $\{\widetilde{g}_{p}\}_{p\in{\mathcal{V}}}$ is the dual frame of $\{g_{p}\}_{p\in{\mathcal{V}}}$ if
@@
In these definitions, the set ${\mathcal{V}}$ can be a cartesian product of both uncountable and countable sets. In particular, countable sets equipped with a counting measure lead to summations instead of integrals. For an in-depth introduction to frames, interested readers can see [2].
@@ 3.1 Fourier domain partitions
A partition of the Fourier domain is defined as a family of disjoint connected open sets $\{\Omega_{n}\}_{n\in\Upsilon}$, with $\Upsilon\in{\mathbb{Z}}$ , of closures covering the Fourier domain, i.e., satisfying
@@
To obtain the partition $\{\Omega_{n}\}_{n\in\Upsilon}$, we can first detect the modes of the Fourier spectrum using the scale-space representation [10] and then define boundaries $\partial\Omega_{n}$ separating these modes. In 1D, the intervals $\Omega_{n}$ can be defined using the lowest minima between these modes. In 2D, the extraction of supports $\Omega_{n}$ can be performed by various methods, such as the Tensor (grid) [6], Ridgelet (radial), Curvelet (radial and angular) [11], Watershed [16] or Voronoi [8] tilings.
A special case is raised by the real-valued functions, since they have a symmetric Fourier spectrum. It is therefore natural to consider a symmetric partition $\{\Omega_{n}\}_{n\in\Upsilon}$, defined as follows.
@@ 3.2 Empirical wavelet filterbank
In this section, we introduce empirical wavelet filter banks induced by a mother wavelet for a given Fourier domain partition. Two types of filters are proposed, depending on the symmetry of the Fourier domain supports.
Let $\psi\in{\mathrm{L}^{2}}({\mathbb{R}}^{N})$ be a mother wavelet such that its Fourier transform $\widehat{\psi}$ is localized in frequency and verifies, for some connected open bounded subset $\Lambda\subseteq\mathrm{supp}\,\psi$,
@@
Given a partition $\{\Omega_{n}\}_{n\in\Upsilon}$, the goal of this section is to define, from any mother wavelet $\psi$, two banks of wavelet filters that are mostly supported in the Fourier domain by $(i)$ $\Omega_{n}$, or $(ii)$ $\Omega_{n}\cup\Omega_{-n}$ in the case of a symmetric partition. To this end, we make the following assumption, which is used throughout the paper.
@@
For every $n\in\Upsilon$, there exists a diffeomorphism $\varphi_{n}$ on ${\mathbb{R}}^{N}$ such that
$\begin{cases}\displaystyle\Lambda=\varphi_{n}(\Omega_{n})&\textrm{if }\Omega_{n}\textrm{ is bounded},\\ \displaystyle\Lambda\subsetneq\varphi_{n}(\Omega_{n})&\textrm{otherwise}.\end{cases}$
A diffeomorphism on a bounded set $\Omega_{n}$ is illustrated in Figure 2. This assumption is mild since, by the Hadamard-Cacciopoli theorem [4], if $\Lambda$ is simply connected (i.e., has no hole), an infinitely differentiable function $\varphi_{n}$ is a diffeomorphism if and only if it is proper, i.e., the preimage of any compact set under $\varphi_{n}$ is compact.
@@
$\widehat{\psi}_{n}(\xi)=\sqrt{|\mathrm{det}\;J_{\varphi_{n}}(\xi)|}\;\widehat{\psi}(\varphi_{n}(\xi)).$
The determinant is a normalization coefficient for the conservation of ernergy when $\Lambda=\varphi_{n}(\Omega_{n})$, i.e.,
@@
We consider the 1D case ($N=1$). Let $\psi$ a mother wavelet on ${\mathbb{R}}^{N}$ of Fourier transform $\widehat{\psi}$ with support $\Lambda=(-\frac{1}{2},\frac{1}{2})$ and $\{\Omega_{n}\}_{n\in\Upsilon}$ a family of non-overlapping open bounded intervals with center $\omega_{n}$ such that ${\mathbb{R}}=\bigcup_{n\in\Upsilon}\overline{\Omega_{n}}$. Then $\varphi_{n}:\xi\mapsto\frac{1}{|\Omega_{n}|}(\xi-\omega_{n})$ is a diffeomorphism such that $\Lambda=\varphi_{n}(\Omega_{n})$ and the empirical wavelet system is given by, for every $\xi\in{\mathbb{R}}$,
@@
This definition is in agreement with the definition given in [7]. For the diffeomorphism $\varphi_{n}:\xi\mapsto\frac{1}{|\Omega_{n}|}(\omega_{n}-\xi)$ also verifying $\Lambda=\varphi_{n}(\Omega_{n})$, the definition becomes, for every $\xi\in{\mathbb{R}}$,
@@
which is different from Equation 1 if $\widehat{\psi}$ is not symmetric (i.e., if $\psi$ is not a real function).
Now, we consider that $\{\Omega_{n}\}_{n\in\Upsilon}$ is a symmetric partition. To build wavelet filters $\phi_{n}$ which are mostly supported by $\Omega_{n}\cup\Omega_{-n}$, we assume that $\Lambda$ satisfies
@@
Necessarily, the system $\{\phi_{n}\}_{n\in\Upsilon}$ must be symmetric with respect to frequency band, that is, it must satisfy the property
$\widehat{\phi}_{n}=\widehat{\phi}_{-n}.$
In this context, we only consider the diffeomorphisms $\varphi_{n}$ for $n\geq 0$. The function $\varphi_{-n}:\xi\mapsto-\varphi_{n}(-\xi)$ is a diffeomorphism, that verifies $\varphi_{-n}(\Omega_{-n})=\Lambda$ when $\varphi_{n}(\Omega_{n})=\Lambda$, which suggests the following definition.
@@
The symmetric empirical wavelet system, denoted $\{\phi_{n}\}_{n\in\Upsilon}$, corresponding to the symmetric partition $\{\Omega_{n}\}_{n\in\Upsilon}$ is defined by, for every $\xi\in{\mathbb{R}}^{N}$,
$\begin{cases}&\displaystyle\widehat{\phi}_{0}(\xi)=\widehat{\psi}_{0}(\xi),\\ &\displaystyle\widehat{\phi}_{n}(\xi)=\frac{1}{\sqrt{2}}\left(\widehat{\psi}_{n}(\xi)+\widehat{\psi}_{-n}(\xi)\right)\quad\textrm{ for }n\neq 0,\end{cases}$ (2)
where $\widehat{\psi}_{n}=\widehat{\psi}\circ\varphi_{n}$ and $\varphi_{-n}:\xi\mapsto-\varphi_{n}(-\xi)$.
This definition does not force the symmetry of $\widehat{\phi}_{n}$ for any $n\in\Upsilon$, but ensures that a symmetry of $\widehat{\psi}_{n}$ implies a symmetry of $\widehat{\phi}_{n}$.
Moreover, we can write Equation 2 explicitely as follows, for every $n\in\Upsilon\setminus\{0\}$ and $\xi\in{\mathbb{R}}^{N}$,
$\widehat{\phi}_{n}(\xi)=\frac{1}{\sqrt{2}}\left(\sqrt{|\mathrm{det}\;J_{\varphi_{n}}(\xi)|}\;\widehat{\psi}(\varphi_{n}(\xi))+\sqrt{|\mathrm{det}\;J_{\varphi_{n}}(-\xi)|}\;\widehat{\psi}(-\varphi_{n}(-\xi))\right).$
Thus, for $n\in\Upsilon\setminus\{0\}$ such that $\widehat{\psi}(\varphi_{n}(\xi))\widehat{\psi}(-\varphi_{n}(-\xi))=0$ for every $\xi\in\Omega_{n}\cup\Omega_{-n}$ and $\Lambda=\varphi_{n}(\Omega_{n})$, the factor $1/\sqrt{2}$ guarantees the conservation of the energy, i.e.,
$\int_{\Omega_{n}\cup\Omega_{-n}}\left|\widehat{\phi_{n}}(\xi)\right|^{2}\mathrm{d}\xi=\int_{\Lambda}\left|\widehat{\psi}(u)\right|^{2}\mathrm{d}u.$
@@
As in Example 3.1, we consider $N=1$, $\widehat{\psi}$ on ${\mathbb{R}}$ with support $\Lambda=(-\frac{1}{2},\frac{1}{2})$ and $\{\Omega_{n}\}_{n\in\Upsilon}$ a family of non-overlapping bounded intervals with center $\omega_{n}$ such that ${\mathbb{R}}=\bigcup_{n\in\Upsilon}\overline{\Omega_{n}}$. For the diffeomorphism $\varphi_{n}:\xi\mapsto\frac{1}{|\Omega_{n}|}(\xi-\omega_{n})$, the symmetric empirical wavelet system reads, for every $\xi\in{\mathbb{R}}$,
$\begin{cases}&\displaystyle\widehat{\phi}_{0}(\xi)=\frac{1}{\sqrt{|\Omega_{0}|}}\widehat{\psi}\left(\frac{\xi}{|\Omega_{0}|}\right),\\ &\displaystyle\widehat{\phi}_{n}(\xi)=\frac{1}{\sqrt{2|\Omega_{n}|}}\left[\widehat{\psi}\left(\frac{\xi-\omega_{n}}{|\Omega_{n}|}\right)+\widehat{\psi}\left(\frac{\xi+\omega_{n}}{|\Omega_{n}|}\right)\right]\quad\textrm{ for }n\neq 0.\end{cases}$
@@
Due to the linearity of the inverse Fourier tranform, the symmetric empirical wavelet system $\{\phi_{n}\}_{n\in\Upsilon}$ also satisfies, in the spatial domain, for every $x\in{\mathbb{R}}^{N}$,
$\phi_{0}(x)=\psi_{0}(x)\quad\textrm{ and }\quad\phi_{n}(x)=\frac{1}{\sqrt{2}}\left(\psi_{n}(x)+\psi_{-n}(x)\right)\quad\textrm{ for }n\neq 0.$ (3)
It is therefore sufficient to write $\psi_{n}$ in the spatial domain to write $\phi_{n}$ in the spatial domain.

### v2

@@ Abstract
The empirical wavelet transform is a data-driven time-scale representation consisting of an adaptive filter bank. Its robustness to data has made it the subject of intense developments and an increasing number of applications in the last decade. However, it has been mostly studied theoretically for signals and its extension to images is limited to a particular generating function. This work presents a general framework for multidimensional empirical wavelet transform based on any wavelet kernel. It also provides conditions to build wavelet frames for both continuous and discrete transforms. Moreover, numerical simulations of transforms are given.
@@ 1 Introduction
The wavelet transform is a reference tool for time-scale representation used in many signal and image processing techniques, such as denoising, deconvolution and texture segmentation. Originally, it consists of projecting data onto wavelet filters that are built from a mother wavelet which is scaled and modulated independently of the data. In practice, this leads to the construction of wavelet filters based on a prescribed scheme, such as a dyadic decomposition. Although this approach is widely used in contemporary research, it is not guaranteed to be optimal for the data at hand, since a prescribed scheme does not take into account the specificity of the underlying Fourier spectrum. Therefore, data-driven filtering approaches have received much attention to provide an accurate time-scale representation that is robust to the data. Among them, empirical mode decomposition [14], a purely algorithmic approach which behaves as a filter bank [6], has been widely used in real-world applications. Inspired by this decomposition, empirical wavelet transform has been introduced in [7] to provide a more consistant decomposition with a sound theoretical foundation [18, 30]. In the last decade, this transform has been the subject of an increasing interest through a continuous development and numerous applications in various fields, as reviewed in [15, 23]. To name a few applications, we can mention seismic time-frequency analysis [22], electrocardiogram (ECG) signal compression [19], epileptic seizure detection [1, 27], speech recognition [21], time series forecasting [25], glaucoma detection [24], hyperspectral image classification [26], texture segmentation [16], multimodal medical image fusion [28] and cancer histopathological image classification [5]. Particularly, this transform has shown to outperform traditional wavelet transforms for texture segmentation [15] and framelet tranforms for texture denoising [17].
@@
The construction of empirical wavelet systems consists of two steps: $(i)$ extracting supports of the harmonic modes of the function under study, and $(ii)$ constructing empirical wavelet filters that are compactly (or very rapidly decaying) supported in the Fourier domain by the extracted supports. This construction is the core of an ongoing and active research effort for one-dimensional (1D) and two-dimensional (2D) functions. For 1D functions, the detection of the segments supporting each harmonic modes in the Fourier domain is usually performed by extracting the lowest minima between them using a scale-space representation [11]. For 2D functions, several different techniques have been proposed to delimit the supports of the harmonic modes, such as the Curvelet [12], Watershed [17] and Voronoi [9] partitioning methods. The construction of wavelet filters based on the detected supports is usually done in the Fourier domain. In the 1D domain, [8] proposed a general framework to build continuous empirical wavelet filters from a generating function, such as the Meyer, Shannon or Gabor scaling functions. These 1D empirical wavelet systems are written as modulations and dilations of a wavelet kernel based on segments that divide the Fourier line. Such systems have been shown to induce both continuous and discrete frames in [10]. In the 2D domain, empirical wavelet filters have been designed following the Littlewood-Palley wavelet formulation for various Fourier partitioning methods [9, 12, 17]. However, the proposed construction is only valid for discrete images and is entirely based on the distance to the support boundaries, which limits its extension to classic scaling functions. So far, no construction of empirical wavelets in higher dimensions has been proposed. In addition, empirical wavelet filters for real-valued functions are built from supports that take into account the symmetry of the corresponding Fourier transform, but these have been little studied theoretically.
@@
This work aims to provide a general framework for empirical wavelet transforms of multidimensional functions, thus extending the 1D framework in [8, 10]. We show that we can build empirical wavelets on Fourier supports, symmetric or not, from any wavelet kernel defined in the Fourier domain, using diffeomorphisms that map these supports to the wavelet kernel’s Fourier support. Both continuous and discrete transforms are considered. In addition, conditions for the construction of wavelet frames are examined.
@@
The paper is organized as follows. The construction of the multidimensional empirical wavelet systems and the resulting transforms is described in Section 3. Theoretical results on these systems as frames are given in Section 4. The special case of Fourier supports resulting from affine deformations of the wavelet kernel’s Fourier support is studied in Section 5. Finally, in Section 6, specific 2D wavelet systems are tailored from classic wavelet kernels and studied numerically on images. A Matlab toolbox will be made publicly available at the time of publication.
@@ 2 Notations and reminders
We respectively denote $\partial\Omega$ and $\overline{\Omega}$ the boundary and closure of a set $\Omega\in{\mathbb{R}}^{N}$. We denote $\Upsilon^{+}=\{n\in\Upsilon\mid n\geq 0\}$ the subset of positive elements of a set $\Upsilon\in{\mathbb{Z}}$. The Jacobian of a differentiable function $\gamma$ is denoted $J_{\gamma}$. We recall that a function $\gamma$ is called a diffeomorphism if it is infinitely differentiable and invertible of inverse infinitely differentiable.
@@
with $({\mathcal{V}},\mu)$ a measure set. In particular, $\{g_{p}\}_{p\in{\mathcal{V}}}$ is called a tight frame if $A_{1}=A_{2}$ and a Parseval frame if $A_{1}=A_{2}=1$. A frame $\{\widetilde{g}_{p}\}_{p\in{\mathcal{V}}}$ is the dual frame of $\{g_{p}\}_{p\in{\mathcal{V}}}$ if, for every $f\in{\mathrm{L}^{2}}({\mathbb{R}}^{N})$,
@@
It is well known that a frame $\{g_{p}\}_{p\in{\mathcal{V}}}$ is a tight frame if and only if there exist $A>0$ such that $\{g_{p}/A\}_{p\in{\mathcal{V}}}$ is a dual frame of $\{g_{p}\}_{p\in{\mathcal{V}}}$. In these definitions, the set ${\mathcal{V}}$ can be a cartesian product of both uncountable and countable sets. In particular, countable sets equipped with a counting measure lead to summations instead of integrals. For an in-depth introduction to frames, interested readers can see [2].
@@ 3.1 Fourier domain partitions
A partition of the Fourier domain is defined as a family of disjoint connected open sets $\{\Omega_{n}\}_{n\in\Upsilon}$, with $\Upsilon\subset{\mathbb{Z}}$, of closures covering the Fourier domain, i.e., satisfying
@@
The partition $\{\Omega_{n}\}_{n\in\Upsilon}$ is usually constructed such that each set $\Omega_{n}$ contains one mode of the Fourier domain. To this end, we can first detect the modes of the Fourier spectrum using the scale-space representation [11] and then define boundaries $\partial\Omega_{n}$ separating these modes. In 1D, the intervals $\Omega_{n}$ can be defined using the lowest minima between these modes. In 2D, the extraction of supports $\Omega_{n}$ can be performed by various methods, such as the Tensor (grid) [7], Ridgelet (radial), Curvelet (radial and angular) [12], Watershed [17] or Voronoi [9] tilings.
A special case is raised by the real-valued functions since their Fourier spectrum is even. It is therefore natural to consider a symmetric partition $\{\Omega_{n}\}_{n\in\Upsilon}$, defined as follows.
@@ 3.2 Empirical wavelet filter bank
In this section, we introduce empirical wavelet filter banks induced by a wavelet kernel for a given Fourier domain partition. Two types of filters are proposed, depending on the symmetry of the Fourier domain supports.
Let $\psi\in{\mathrm{L}^{2}}({\mathbb{R}}^{N})$ be a wavelet kernel such that its Fourier transform $\widehat{\psi}$ is localized in frequency and verifies, for some connected open bounded subset $\Lambda\subseteq\mathrm{supp}\,\psi$,
@@
Given a partition $\{\Omega_{n}\}_{n\in\Upsilon}$, the goal of this section is to define, from any wavelet kernel $\psi$, two banks of wavelet filters that are mostly supported in the Fourier domain by $(i)$ $\Omega_{n}$, or $(ii)$ $\Omega_{n}\cup\Omega_{-n}$ in the case of a symmetric partition. To this end, we make the following assumption, which is used throughout the paper.
@@
For every $n\in\Upsilon$, there exists a diffeomorphism $\gamma_{n}$ on ${\mathbb{R}}^{N}$ such that
$\begin{cases}\displaystyle\Lambda=\gamma_{n}(\Omega_{n})&\textrm{if }\Omega_{n}\textrm{ is bounded},\\ \displaystyle\Lambda\subsetneq\gamma_{n}(\Omega_{n})&\textrm{otherwise}.\end{cases}$
A diffeomorphism on a bounded set $\Omega_{n}$ is illustrated in Figure 2. This assumption is mild since, by the Hadamard-Cacciopoli theorem [4], if $\Lambda$ is simply connected (i.e., has no hole), an infinitely differentiable function $\gamma_{n}$ is a diffeomorphism if and only if it is proper, i.e., the preimage of any compact set under $\gamma_{n}$ is compact.
@@
$\widehat{\psi}_{n}(\xi)=\sqrt{|\mathrm{det}\;J_{\gamma_{n}}(\xi)|}\;\widehat{\psi}(\gamma_{n}(\xi)).$
The determinant is a normalization coefficient for the conservation of energy when $\Lambda=\gamma_{n}(\Omega_{n})$, i.e.,
@@
We consider the 1D case ($N=1$). Let $\psi$ a wavelet kernel on ${\mathbb{R}}$ of Fourier transform $\widehat{\psi}$ with support $\Lambda=(-\frac{1}{2},\frac{1}{2})$ and $\{\Omega_{n}\}_{n\in\Upsilon}$ a family of non-overlapping open bounded intervals with center $\omega_{n}$ such that ${\mathbb{R}}=\bigcup_{n\in\Upsilon}\overline{\Omega_{n}}$. Then $\gamma_{n}:\xi\mapsto\frac{1}{|\Omega_{n}|}(\xi-\omega_{n})$ are diffeomorphisms such that $\Lambda=\gamma_{n}(\Omega_{n})$ and the empirical wavelet system is given by, for every $\xi\in{\mathbb{R}}$,
@@
This definition is in agreement with the definition given in [8]. For the diffeomorphisms $\kappa_{n}:\xi\mapsto\frac{1}{|\Omega_{n}|}(\omega_{n}-\xi)$ also satisfying $\Lambda=\kappa_{n}(\Omega_{n})$, Definition 1 becomes, for every $\xi\in{\mathbb{R}}$,
@@
which is different from Equation 1 if $\widehat{\psi}$ is not even.
Now, we consider that $\{\Omega_{n}\}_{n\in\Upsilon}$ is a symmetric partition. To build wavelet filters $\chi_{n}$ which are mostly supported by $\Omega_{n}\cup\Omega_{-n}$, we assume that $\Lambda$ satisfies
@@
Necessarily, the system $\{\chi_{n}\}_{n\in\Upsilon}$ must be symmetric with respect to frequency band, that is, it must satisfy the property
$\widehat{\chi}_{n}=\widehat{\chi}_{-n}.$
In this context, we only consider the diffeomorphisms $\gamma_{n}$ for $n\geq 0$. The function $\gamma_{-n}:\xi\mapsto-\gamma_{n}(-\xi)$ is a diffeomorphism, that verifies $\gamma_{-n}(\Omega_{-n})=\Lambda$ when $\gamma_{n}(\Omega_{n})=\Lambda$, which suggests the following definition.
@@
The symmetric empirical wavelet system, denoted $\{\chi_{n}\}_{n\in\Upsilon}$, corresponding to the symmetric partition $\{\Omega_{n}\}_{n\in\Upsilon}$ is defined by, for every $\xi\in{\mathbb{R}}^{N}$,
$\begin{cases}&\displaystyle\widehat{\chi}_{0}(\xi)=\widehat{\psi}_{0}(\xi),\\ &\displaystyle\widehat{\chi}_{n}(\xi)=\frac{1}{\sqrt{2}}\left(\widehat{\psi}_{n}(\xi)+\widehat{\psi}_{-n}(\xi)\right)\quad\textrm{ for }n\neq 0,\end{cases}$ (2)
where $\widehat{\psi}_{n}=\widehat{\psi}\circ\gamma_{n}$ and $\gamma_{-n}:\xi\mapsto-\gamma_{n}(-\xi)$.
In this definition, $\widehat{\chi}_{n}$ is not necessarily even for any $n\in\Upsilon$, but ensures that the parity of $\widehat{\psi}_{n}$ implies the parity of $\widehat{\chi}_{n}$. Moreover, we can write Equation 2 explicitely as follows, for every $n\in\Upsilon\setminus\{0\}$ and $\xi\in{\mathbb{R}}^{N}$,
$\widehat{\chi}_{n}(\xi)=\frac{1}{\sqrt{2}}\left(\sqrt{|\mathrm{det}\;J_{\gamma_{n}}(\xi)|}\;\widehat{\psi}(\gamma_{n}(\xi))+\sqrt{|\mathrm{det}\;J_{\gamma_{n}}(-\xi)|}\;\widehat{\psi}(-\gamma_{n}(-\xi))\right).$
Thus, for $n\in\Upsilon\setminus\{0\}$ such that $\int_{\Omega_{n}\cup\Omega_{-n}}\widehat{\psi}(\gamma_{n}(\xi))\overline{\widehat{\psi}(-\gamma_{n}(-\xi))}=0$ and $\Lambda=\gamma_{n}(\Omega_{n})$, the factor $1/\sqrt{2}$ guarantees the conservation of the energy, i.e.,
$\int_{\Omega_{n}\cup\Omega_{-n}}\left|\widehat{\chi}_{n}(\xi)\right|^{2}\mathrm{d}\xi=\int_{\Lambda}\left|\widehat{\psi}(u)\right|^{2}\mathrm{d}u.$
@@
As in Example 3.1, we consider $N=1$, $\widehat{\psi}$ on ${\mathbb{R}}$ with support $\Lambda=(-\frac{1}{2},\frac{1}{2})$ and $\{\Omega_{n}\}_{n\in\Upsilon}$ a family of non-overlapping bounded intervals with center $\omega_{n}$ such that ${\mathbb{R}}=\bigcup_{n\in\Upsilon}\overline{\Omega_{n}}$. For the diffeomorphism $\gamma_{n}:\xi\mapsto\frac{1}{|\Omega_{n}|}(\xi-\omega_{n})$, the symmetric empirical wavelet system reads, for every $\xi\in{\mathbb{R}}$,
$\begin{cases}&\displaystyle\widehat{\chi}_{0}(\xi)=\frac{1}{\sqrt{|\Omega_{0}|}}\widehat{\psi}\left(\frac{\xi}{|\Omega_{0}|}\right),\\ &\displaystyle\widehat{\chi}_{n}(\xi)=\frac{1}{\sqrt{2|\Omega_{n}|}}\left[\widehat{\psi}\left(\frac{\xi-\omega_{n}}{|\Omega_{n}|}\right)+\widehat{\psi}\left(\frac{\xi+\omega_{n}}{|\Omega_{n}|}\right)\right]\quad\textrm{ for }n\neq 0.\end{cases}$
@@
Due to the linearity of the inverse Fourier tranform, the symmetric empirical wavelet system $\{\chi_{n}\}_{n\in\Upsilon}$ also satisfies, in the spatial domain, for every $x\in{\mathbb{R}}^{N}$,
$\chi_{0}(x)=\psi_{0}(x)\quad\textrm{ and }\quad\chi_{n}(x)=\frac{1}{\sqrt{2}}\left(\psi_{n}(x)+\psi_{-n}(x)\right)\quad\textrm{ for }n\neq 0.$ (3)
It is therefore sufficient to write $\psi_{n}$ in the spatial domain to write $\chi_{n}$ in the spatial domain.